OpenStudy (anonymous):

Explain, in complete sentences, how you would use the elimination method to solve the following system of equations. Provide the solution to the system. 6x – 5y = 4 9x + 4y = –17

OpenStudy (anonymous):

multiply first equation by 4 and the second by 5 24x - 20y = 16 45x + 20y = -85 -adding these two equations eliminates y 69x = -69 x = -1 now we can find the value of y by plugging x=-1 into one of the above equations 6(-1) - 5y = 4 -6 - 5y = 4 5y = -10 y = -2 x=-1 y=-2 you can check these results by plugging them into second equaion: 9(-1) + 4(-2) = -17 - checks out